# On Geometry and Topology of Hessian Manifolds

### v4

@@ 5.4 Topology of Hessian 4-Manifolds
If $(M,g)$ is of negative sectional curvature, then, by Hopf’s conjecture in dimension four which is proved using Gauß–Bonnet formula, we have that $\chi(M)>0$, a contradiction to Proposition 2.17. ∎
Proposition 5.22.
Let $((M,\nabla),h)$ be a closed Hessian $4$-manifold. Suppose that $g$ is an Einstein metric on $M$. Then $(M,g)$ is a flat Riemannian manifold.
Proof.
By considering the orientation double cover, we may assume without loss of generality that $M$ is orientable.
Now, assume the contrary that the Einstein manifold $(M,g)$ is not flat. Then, since by Lemma 5.17
$\chi(M)=0=\frac{3}{2}\left|\operatorname{sign}\left(Q_{M}\right)\right|,$
by Hitchin–Thorpe inequality we have that the universal covering of $M$ is $\pi\colon X\rightarrow M$ where $X$ is a K3 surface. Since K3 surfaces are compact and without loss of generality assume $\pi$ is orientation-preserving, we have that $d:=\operatorname{deg}(\pi)<\infty$ is a positive integer. But as $\chi(X)=d\cdot\chi(M)$ by Riemann-Hurwitz formula, and it is well-known in algebraic geometry that $\chi(X)=24$, we arrive at $\chi(M)>0$, which contradicts Proposition 2.17. ∎
Appendix A On Homotopy Tori
Theorem 2.22 crucially uses the result in geometric topology that a homotopy torus is in fact homeomorphic to the flat torus, which we will now produce a proof for completeness. We shall first recall the following two propositions from low-dimensional topology.
Proposition A.1.
Let $M$ be a closed $3$-manifold. Suppose that there exists a closed surface $B$ with infinite fundamental group, and a finitely generated normal subgroup $G$ of $\pi_{1}(M)$, such that the following sequence
$1\longrightarrow G\longrightarrow\pi_{1}(M)\longrightarrow\pi_{1}(B)\longrightarrow 1$
is exact. Then $G=\mathbb{Z}$ and $M$ is an $\mathbb{S}^{1}$-bundle over $B$.
Proof.
See [33] for a combinatorial proof of this statement assuming Poincaré’s conjecture, while the latter is confirmed by G.Perelman in [34],[35], and [36]. See also [37]. ∎
Proposition A.2.
Let $f\colon M\rightarrow N$ be a homotopy equivalence between two closed aspherical 4-manifolds with polycyclic fundamental group. Then $f$ is homotopic to a homeomorphism.
Proof.
This is proved in [38]. ∎
Now we are ready to prove the main theorem regarding homotopy tori.
Theorem A.3.
Let $M$ be a closed topological manifold. If $M$ is homotopy equivalent to the flat $n$-torus $\mathbb{T}^{n}$, then $M$ is actually homeomorphic to $\mathbb{T}^{n}$.
Proof.
Comparing the top homology groups yields that $M$ is orientable and $\operatorname{dim}M=n$. For $n\leq 2$, it is well-known that closed $n$-manifolds are diffeomorphic if they are of the same homotopy type. For $n\geq 5$, the desired result is proved in [39]. It remains to consider the case when $n=3$ or $4$.
Assume that $n=4$. Since $\mathbb{Z}^{4}$ is polycyclic and the closed 4-manifold $\mathbb{T}^{4}$ is aspherical, by Proposition A.2 we conclude that $M$ is homeomorphic to $\mathbb{T}^{4}$.
Assume that $n=3$. By Proposition A.1, it is readily seen that $M$ is a $\mathbb{S}^{1}$-bundle over $\mathbb{T}^{2}$. But by the theory of 3-manifolds, two $\mathbb{S}^{1}$-bundles over $\mathbb{T}^{2}$ are homeomorphic if and only if their Euler class coincide. Assume that $M$ is the $\mathbb{S}^{1}$-bundles over $\mathbb{T}^{2}$ with Euler class $k\in\mathbb{Z}$. Then $\mathrm{H}_{1}(M;\mathbb{Z})=\mathbb{Z}^{2}\oplus\mathbb{Z}/k\mathbb{Z}$. But by Hurewicz’s theorem, $\mathrm{H}_{1}(M;\mathbb{Z})=\mathbb{Z}^{3}$ as the Abelianization of $\pi_{1}(M)$. Therefore we obtain that $k=0$, and hence $M$ is the trivial $\mathbb{S}^{1}$-bundles over $\mathbb{T}^{2}$, i.e. $M$ is homeomorphic to the 3-torus $\mathbb{T}^{3}$. ∎
Appendix B Pushforward Parallel Vectors along Harmonic Map
In this section, we prove that the length square of the pushforward of a parallel vector field under a harmonic mapping into a negatively curved space is subharmonic, which is very crucially used in the proof of Theorem 4.10.
We first introduce a preparatory lemma.
Lemma B.1.
Let $(M,g)$ and $(N,h)$ be Riemannian manifolds. Let $f\colon M\rightarrow N$ be a smooth mapping. Let $\hat{\nabla}:=\operatorname{Id}\otimes f^{*}\nabla^{h}+\nabla^{g}\otimes\operatorname{Id}$ be the induced connection on the tensor bundle $\mathscr{E}:=T^{*}M\otimes f^{*}TN$, where $\nabla^{g}$ is the Levi-Civita connection of $(M,g)$ and $\nabla^{h}$ is the Levi-Civita connection of $(N,h)$. Then, for any vector fields $X,Y,Z$ on $M$ and smooth section $\alpha$ of $\mathscr{E}\rightarrow M$, it holds that
$(\hat{\nabla}_{X}\alpha)(Y)=(f^{*}\nabla^{h})_{X}(\alpha(Y))-\alpha(\nabla^{g}_{X}Y)$
and that
$(\hat{\nabla}_{X}\hat{\nabla}_{Y}\alpha)(Z)=$ $(f^{*}\nabla^{h})_{X}(f^{*}\nabla^{h})_{Y}(\alpha(Z))-(f^{*}\nabla^{h})_{X}(\alpha(\nabla^{g}_{Y}Z))$
$-(f^{*}\nabla^{h})_{Y}(\alpha(\nabla^{g}_{X}Z))+\alpha(\nabla^{g}_{Y}\nabla^{g}_{X}Z).$
Proof.
The first identity follows directly from Leibniz rule and the definition of $\hat{\nabla}$.
Repeatedly applying the first identity yields
$(\hat{\nabla}_{X}\hat{\nabla}_{Y}\alpha)(Z)=$ $(f^{*}\nabla^{h})_{X}((\hat{\nabla}_{Y}\alpha)(Z))-(\hat{\nabla}_{Y}\alpha)(\nabla^{g}_{X}Z)$
$=$ $(f^{*}\nabla^{h})_{X}((f^{*}\nabla^{h})_{Y}(\alpha(Z))-\alpha(\nabla^{g}_{Y}Z))$
$-((f^{*}\nabla^{h})_{Y}(\alpha(\nabla^{g}_{X}Z))-\alpha(\nabla^{g}_{Y}\nabla^{g}_{X}Z))$
$=$ $(f^{*}\nabla^{h})_{X}(f^{*}\nabla^{h})_{Y}(\alpha(Z))-(f^{*}\nabla^{h})_{X}(\alpha(\nabla^{g}_{Y}Z))$
$-(f^{*}\nabla^{h})_{Y}(\alpha(\nabla^{g}_{X}Z))+\alpha(\nabla^{g}_{Y}\nabla^{g}_{X}Z)$
as desired. ∎
The following theorem as it stated is not only a tool to prove Theorem 4.10, but also reveals an important property of harmonic mappings.
Theorem B.2.
Let $(M,g)$ and $(N,h)$ be Riemannian manifolds with $\operatorname{dim}(N)\geq 2$. Let $f\colon M\rightarrow N$ be a harmonic mapping. Suppose that $(N,h)$ is of non-positive sectional curvature, and that $X$ is a parallel vector field on $(M,g)$. Then the energy density
$\frac{1}{2}(f^{*}h)(f_{*}X,f_{*}X)\in C^{\infty}(M)$
is a subharmonic function on $(M,g)$. Moreover, if we assume further that $M$ is compact and $(N,h)$ is of negative sectional curvature, then $f_{*}X\equiv 0$.
Proof.
Consider orthonormal frame fields $e_{1},\dots,e_{n}$ on $(M,g)$.
Let $\hat{\nabla}:=\operatorname{Id}\otimes f^{*}\nabla^{h}+\nabla^{g}\otimes\operatorname{Id}$ be the induced connection on the tensor bundle $\mathscr{E}:=T^{*}M\otimes f^{*}TN$, where $\nabla^{g}$ is the Levi-Civita connection of $(M,g)$ and $\nabla^{h}$ is the Levi-Civita connection of $(N,h)$.
Since $X$ is a parallel vector field on $(M,g)$ and $\hat{\nabla}(df)$ is symmetric, we have
$(\hat{\nabla}_{e_{i}}\hat{\nabla}_{X}(df))(e_{j})=$ $(f^{*}\nabla^{h})_{e_{i}}((\hat{\nabla}_{X}(df))(e_{j}))-(\hat{\nabla}_{X}(df))(\nabla^{g}_{e_{i}}e_{j})$
$=$ $(f^{*}\nabla^{h})_{e_{i}}((\hat{\nabla}_{e_{j}}(df))(X))-(\hat{\nabla}_{e_{j}}(df))(\nabla^{g}_{e_{i}}X)-(\hat{\nabla}_{X}(df))(\nabla^{g}_{e_{i}}e_{j})$
$=$ $(\hat{\nabla}_{e_{i}}\hat{\nabla}_{e_{j}}(df))(X)-(\hat{\nabla}_{X}(df))(\nabla^{g}_{e_{i}}e_{j})$
by Lemma B.1. Denote by $\hat{\Delta}$ the connection Laplacian of $\hat{\nabla}$, that is,
$(\hat{\Delta}(df))(X)=\sum_{i=1}^{n}(\hat{\nabla}_{e_{i}}\hat{\nabla}_{e_{i}}(df))(X)-\sum_{i=1}^{n}(\hat{\nabla}_{X}(df))(\nabla^{g}_{e_{i}}e_{j}).$
Again, since $X$ is a parallel vector field on $(M,g)$, we obtain $\operatorname{Ric}^{g}(-,X)=0$ and hence
$\sum_{i=1}^{n}R^{g}(X,e_{i})e_{i}=0$
where $R^{g}$ is the Riemann curvature tensor and $\operatorname{Ric}^{g}$ is the Ricci tensor of $(M,g)$.
Therefore, by the product rule for curvature tensors, we have
$\sum_{i=1}^{n}(\hat{R}(X,e_{i})(df))(e_{i})=$ $\sum_{i=1}^{n}R^{h}(f_{*}X,f_{*}e_{i})(f_{*}e_{i})-f_{*}\sum_{i=1}^{n}R^{g}(X,e_{i})e_{i}$
$=$ $\sum_{i=1}^{n}R^{h}(f_{*}X,f_{*}e_{i})(f_{*}e_{i})$
where $R^{h}$ is the curvature tensor of $\nabla^{h}$, and $\hat{R}$ is the curvature tensor of $\hat{\nabla}$.
Now, since $f$ is harmonic, a computation yields
$0=$ $\sum_{i=1}^{n}(f^{*}\nabla^{h})_{X}((\hat{\nabla}_{e_{i}}(df))(e_{i}))$
$=$ $\sum_{i=1}^{n}(\hat{\nabla}_{e_{i}}\hat{\nabla}_{X}(df))(e_{i})+\sum_{i=1}^{n}(\hat{R}(X,e_{i})(df))(e_{i})$
$=$ $(\hat{\Delta}(df))(X)+\sum_{i=1}^{n}R^{h}(f_{*}X,f_{*}e_{i})(f_{*}e_{i}).$
For simplicity, for any sections $s,s_{1},s_{2}$ of $f^{*}TN$, we write $|s|:=(f^{*}h)(s,s)$ and
$\langle s_{1},s_{2}\rangle:=(f^{*}h)(s_{1},s_{2}).$
Since $(N,h)$ is of non-positive sectional curvature, we have
$\langle(\hat{\Delta}(df))(X),f_{*}X\rangle=\sum_{i=1}^{n}\langle R^{h}(f_{*}e_{i},f_{*}X)(f_{*}e_{i}),f_{*}X\rangle\geq 0.$
Since $X$ is a parallel vector field on $(M,g)$, Weitzenböck identity reads
$\frac{1}{2}\Delta_{g}|f_{*}X|^{2}=\langle(\hat{\Delta}(df))(X),f_{*}X\rangle+\sum_{i=1}^{n}|(f^{*}\nabla^{h})_{e_{i}}(f_{*}X)|^{2}\geq 0$
where $\Delta_{g}$ here is the Laplace-Beltrami operator on $(M,g)$. The proof of the first claim is completed.
Now, we assume further that $M$ is compact. Then, by Liouville’s theorem, the energy density
$u:=\frac{1}{2}(f^{*}h)(f_{*}X,f_{*}X)\in C^{\infty}(M)$
is a constant function. Since $\Delta_{g}u=0$ and $\langle(\hat{\Delta}(df))(X),f_{*}X\rangle\geq 0$, we obtain
$\frac{1}{2}\Delta_{g}|f_{*}X|^{2}=\langle(\hat{\Delta}(df))(X),f_{*}X\rangle=\sum_{i=1}^{n}|(f^{*}\nabla^{h})_{e_{i}}(f_{*}X)|^{2}=0$
and hence
$\sum_{i=1}^{n}\langle R^{h}(f_{*}e_{i},f_{*}X)(f_{*}e_{i}),f_{*}X\rangle=0.$
Since $\langle R^{h}(f_{*}e_{i},f_{*}X)(f_{*}e_{i}),f_{*}X\rangle\geq 0$, we also have that
$\langle R^{h}(f_{*}e_{i},f_{*}X)(f_{*}e_{i}),f_{*}X\rangle=0$
for all $i=1,\dots,n$. Therefore, in particular, for any smooth vector field $Y$ on $N$, we have $\langle R^{h}(Y,f_{*}X)(Y),f_{*}X\rangle=0$. If $f_{*}X$ is non-vanishing at some point $p\in M$, then by the additional assumption on sectional curvatures, for any vector $\xi\in T_{f(p)}N$ that is not proportional to $f_{*}X$ at $f(p)\in N$, we have
$\langle R^{h}(\xi,f_{*}X)(\xi),f_{*}X\rangle>0$
which is a contradiction. Therefore we conclude that $f_{*}X\equiv 0$. ∎
Acknowledgement
The author wants to express the author’s most sincere gratitude to Weiyi Zhang for very useful discussions and general advices along the whole preparation process of this manuscript, and to Maxwell Stolarski for a careful proof reading with various valuable suggestions. The author also wants to thank Jun Zhang for pointing out a potential connection with Riemann extension, which results in an improvement of section3.2.
Special thanks should also go to Saul Schleimer who inspires the author’s work on Hessian 3-folds, and to Ruadhaí Dervan for a discussion about Kähler geometry of Hessian manifolds.
@@ Acknowledgement
The author is in debt to the Russian mathematical society, especially, to all the professors teaching the math in Moscow program, and most importantly, to the author’s supervisor Alexander Petrovich Veselov at Loughborough University, as they cultivated the author’s mathematical literacy and maturity.
Finally, the author would love to thank the author’s family and friends, for their kindness and steady help.
Statements and Declarations
No funding was received to assist with the preparation of this manuscript.
The author certifies that the author has no affiliations with or involvement in any other organization or entity with any financial interest or non-financial interest in the subject matter or materials discussed in this manuscript.
This article is licensed under a Creative Commons Attribution 4.0 International License, which permits use, sharing, adaptation, distribution and reproduction in any medium or format, as long as you give appropriate credit to the original author and the source, provide a link to the Creative Commons licence, and indicate if changes were made. The images or other third party material in this article are included in the article’s Creative Commons licence, unless indicated otherwise in a credit line to the material. If material is not included in the article’s Creative Commons licence and your intended use is not permitted by statutory regulation or exceeds the permitted use, you will need to obtain permission directly from the copyright holder.
The author hereby provide consent for the publication of the manuscript detailed above.

### v6

@@ 5.3 Closed Orientable Hessian 4-Manifolds
By Theorem 4.15, there exists a finite-sheet covering map $\pi\colon E\rightarrow M$ such that the lifting $(E,\pi^{*}\nabla)$ is the product of a compact hyperbolic affine manifold with a flat torus of dimension $d\in\{0,\dots,4\}$.
If $d\geq 3$, then the desired result holds trivially.
If $1\leq d\leq 2$, then $M$ is diffeomorphic to a Bieberbach 4-manifold or a $\Sigma_{g}$-bundle over $\mathbb{T}^{2}$ with periodic monodromy, by Theorem 5.6 and Theorem 5.9.
Now, suppose that $d=0$. Then $(M,\nabla)$ itself is a compact orientable hyperbolic affine manifold. By Corollary 4.12, there exists a fiber bundle $f\colon M\rightarrow\mathbb{S}^{1}$ of which fiber is a closed orientable $3$-manifold NN, such that the monodromy $\phi]\in\operatorname{Mod}(N)$ of the mapping torus $f\colon M\rightarrow\mathbb{S}^{1}$ is periodic. Since $N$ is aspherical as $M$ is, in particular the second homotopy group $\pi_{2}(N)=0$, and hence $N$ is irreducible. If $N$ is atoroidal, then by the hyperbolization theorem, there exists a Riemannian metric $g$ on $N$ such that $(N,g)$ is an hyperbolic $3$-manifold, and that $\phi$ is isotopic to an isometry of $(N,g)$. From now on, we shall assume that $N$ is not atoroidal. Then by Thurston’s geometrization, we have that $N$ is a graph manifold. Therefore, the fundamental group $\pi_{1}(M)$ admits a solvable subgroup of finite index. In view of Lemma 5.11 and Proposition 5.12, the Enriques–Kodaira classification yields that $M$ is diffeomorphic to one of the following complex surfaces:
  1. 1.
    a complex 2-torus;
  2. 2.
    a primary Kodaira surface;
  3. 3.
    a hyperelliptic surface;
  4. 4.
    an Inoue surface,
while complex 2-tori and hyperelliptic surfaces are all diffeomorphic to Bieberbach 4-manifolds, and the primary Kodaira surfaces and Inoue surfaces are not mapping tori of any periodic mapping classes.
The desired result is therefore proved. ∎
Remark 5.16.
Analogously to Example 5.7 and Example 5.8, examples for Theorem 5.15 can all be produced. Here we omit the detailed construction and leave it to the cautious readers.
Acknowledgements
The author wants to express the author’s most sincere gratitude to Weiyi Zhang for very useful discussions and general advice along the whole preparation process of this manuscript, and to Maxwell Stolarski for a careful proofreading with various valuable suggestions. Special thanks should also go to Saul Schleimer who inspires the author’s work on Hessian 3-manifolds.